You have two square plots formed with 100m of fencing Determine the equation of the curve of best fit, and use the equation to determine the value of x that minimizes the area of each square.
Answer by stanbon(75887)

You have two square plots formed with 100m of fencing Determine the equation of the curve of best fit, and use the equation to determine the value of x that minimizes the area of each square.
-------------
Perimeter of one of the squares:: x
side = x/4
Area = (x/4)^2 = x^2/16
A'(x) = (1/16)(2x) = (1/8)x
Solve:: (1/16)x = 0
x = 0
---------------------
Perimeter of other square:: 100-x
side = (100-x)/4
Area = [(100-x)/4]^2 = (100-x)^2/16
A'(x) = (1/16)(2(100-x)(-1) = (-1/8)(100-x)
Solve: 100-x = 0
x = 100
----------------------
Cheers,
Stan H.